Two music CDs and four music cassette tapes cost a total of $40.00. However, three music CDs and five cassette tapes cost $55.00. Find the price of each.
So lost with this problem, please help.
Thanks.

Answer by lwsshak3(11628)

Two music CDs and four music cassette tapes cost a total of $40.00. However, three music CDs and five cassette tapes cost $55.00. Find the price of each.
***
let x=price of CD's
let y=price of cassette tapes
..
2x+4y=40
3x+5y=55
..
6x+12y=120 (mult.eq by 3)
6x+10y=110 (mult.eq by 2)
subtract:
2y=10
y=5
2x=40-4y=40-20=20
x=10
price of CD's=$10/each
price of cassette tapes=$5/each